ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 168, 2008

Adan-Bante, E. Derived length and products of conjugacy classes  93

Akiyama, S., C. Frougny and J. Sakarovitch Powers of rationals modulo 1 and rational base number systems

53

Balmuş A., S. Montaldo and C. Oniciuc Classification results for biharmonic submanifolds in spheres

201

Bartholdi, L. On amenability of group algebras, I  153
Berkovich, Y. Hall chains in normal subgroups of finite p-groups  29

Bisztriczky, T., F. Fodor and D. Oliveros The T (4) property of families of unit disks

239

Burguet, D. A proof of Yomdin-Gromov’s algebraic lemma  291

Fodor, F. See T. Bisztriczky Frougny, C. See S. Akiyama

Fukuda, D. Classification of Hopf Algebras of Dimension 18  119

Golasiński, M. and F. Gómez Ruiz On polynomial automorphisms of spheres

275

Gómez Ruiz, F. See M. Golasiński Grünbaum, B., T. Kaiser, D. Král’ and M. Rosenfeld Equipar-

tite graphs

431

Groves, D. and J. F. Manning Dehn filling in relatively hyperbolic groups

317

Hájek, P. and M. Johanis Polyhedrality in Orlicz spaces  167
Ho, Wing Kai Manifolds without \frac{1}{k} -geodesics  189

Johanis, M. See P. Hájek Kaiser, T. See B. Grünbaum Král’, D. See B. Grünbaum Machura, M. and B. Tsaban The combinatorics of the Baer-

Specker group

125

Manning, J. F. See D. Groves Mátrai, T. Resolvent norm decay does not characterize norm

continuity

1

Montaldo, S. See A. Balmuş Oliveros, D. See T. Bisztriczky Oniciuc, C. See A. Balmuş Rosenfeld, M. See B. Grünbaum Ryazanov, V. and E. Sevost’yanov Toward the theory of ring

Q–homeomophisms

101

Sakarovitch, J. See S. Akiyama Sevost’yanov, E. See V. Ryazanov Tsaban, B. See M. Machura

Valdimarsson, S. I. Optimisers for the Brascamp–Lieb inequality  253

Yaskin, V. Modified Shephard’s problem on projections of convex bodies 221